OpenStudy (anonymous):

P= 2l + 2w, solve for l

OpenStudy (zzr0ck3r):

subtract 2w from both sides then divide everything by 2

OpenStudy (jim766):

p = 2L + 2W -2W - 2W p - 2W = 2L ----- -- 2 2

OpenStudy (anonymous):

A= 1/2bh, solve for h

OpenStudy (anonymous):

@Jim766 help

OpenStudy (johnweldon1993):

A = 1/2bh multiply both sides by 2 2A = bh divide both sides by 'b' 2A = bh -------- b b h = 2A --- b